### v2

# Efficient Dual-domain Image Dehazing with Haze Prior Perception

Abstract
Transformers offer strong global modeling for single-image dehazing but come with high computational costs. Most methods rely on spatial features to capture long-range dependencies, making them less effective under complex haze conditions. Although some integrate frequency-domain cues, weak coupling between spatial and frequency branches limits their performance. To address these issues, we propose the Dark Channel Guided Frequency-aware Dehazing Network (DGFDNet), a dual-domain framework that explicitly aligns degradation across spatial and frequency domains. At its core, the DGFDBlock consists of two key modules: 1) Haze-Aware Frequency Modulator (HAFM), which uses dark channel priors to generate a haze confidence map for adaptive frequency modulation, achieving global degradation-aware spectral filtering. 2) Multi-level Gating Aggregation Module (MGAM), which fuses multi-scale features via multi-scale convolutions and a hybrid gating mechanism to recover fine-grained structures. Additionally, the Prior Correction Guidance Branch (PCGB) incorporates feedback for iterative refinement of the prior, improving haze localization accuracy, particularly in outdoor scenes. Extensive experiments on four benchmark datasets demonstrate that DGFDNet achieves state-of-the-art performance with improved robustness and real-time efficiency. Code is available at: https://github.com/Dilizlr/DGFDNet.
I Introduction
Hazy images suffer from contrast degradation, color distortion, and detail loss, posing challenges for downstream tasks such as object detection [1, 2] and semantic segmentation [3, 4]. Single-image dehazing seeks to recover a clear scene from a hazy input, but it is an ill-posed problem due to the spatially varying and unknown haze distribution.
Early methods [5, 6, 7, 8] employ handcrafted priors to estimate haze distribution, effective in simple scenes but underperforming in more complex ones. CNN-based methods [9, 10, 11, 12] improve reconstruction quality through end-to-end learning, but their limited receptive fields hindered global dependency modeling, which is essential for restoring structural consistency and lost details. Transformer-based models [13, 14, 15] excel at capturing long-range dependencies via self-attention mechanisms, advancing dehazing performance. However, their quadratic complexity restricts real-time applications.
Recent work [16, 17, 18, 19] has explored integrating frequency-domain information to improve global context modeling efficiency. This is driven by two key insights: 1) Haze primarily distorts the real and magnitude components of the Fourier spectrum, while the phase and imaginary parts retain structural integrity (Fig. 2 (a)), making frequency learning ideal for capturing haze properties. 2) Local frequency changes induce global spatial effects (Fig. 2 (b)), suggesting that frequency-based processing is more efficient for modeling long-range dependencies. However, existing spatial-frequency fusion methods typically use loosely coupled dual-branch designs, lacking explicit degradation alignment, which limits information exchange and robustness in complex haze conditions.
Figure 1: Comparison results on the SOTS-Indoor dataset. The bubble size represents the number of model parameters, and the number below each model indicates the SSIM value.
To overcome these limitations, we propose the Dark Channel Guided Frequency-aware Dehazing Network (DGFDNet), a novel dual-domain framework that explicitly aligns haze degradation cues across spatial and frequency domains, enhancing robustness and achieving a compact design. By combining spatial haze localization with targeted frequency-domain modulation, DGFDNet effectively captures global context while maintaining computational efficiency, striking a balance between performance and practicality (Fig. 1).
Figure 2: (a) Investigation of haze degradation in the frequency domain by separately swapping phase and imaginary components between hazy and clean images.(b) Demonstration of the global impact of frequency-domain operations on spatial domain by modifying local amplitude regions in hazy images.
DGFDNet consists of multi-scale DGFDBlocks, each containing two key modules: the Haze-Aware Frequency Modulator (HAFM) and the Multi-level Gating Aggregation Module (MGAM). Guided by the physically grounded dark channel prior, HAFM generates a pixel-level haze confidence map and selectively modulates frequency components related to haze, enabling global degradation modeling. Building on this, MGAM refines spatial representations through multi-scale convolutions and a hybrid gating mechanism, enhancing fine-grained perception and recovering high-frequency details.
While the dark channel prior provides a coarse haze estimate, it struggles in complex outdoor scenes. To tackle this, we introduce the Prior Correction Guided Branch (PCGB), which not only delivers haze degradation cues to HAFM but also receives refinement feedback from MGAM, iteratively correcting the prior information. This closed-loop feedback mechanism enforces consistent alignment of multi-stage, multi-domain degradation information, improving haze localization accuracy and robustness in challenging real-world conditions.
Each DGFDBlock jointly optimizes spatial and frequency-domain features under PCGB guidance, enabling efficient global context modeling while preserving local details. Our main contributions are summarized as follows:
1) We propose DGFDNet, a dual-domain dehazing framework that explicitly aligns spatial and frequency degradation cues. Each DGFDBlock integrates HAFM for haze-aware spectral modulation and MGAM for detail enhancement, balancing global context modeling and local restoration.
2) We design PCGB, which employs a closed-loop feedback mechanism to iteratively refine the dark channel guidance. This dynamic correction strategy significantly improves haze localization and model robustness in complex outdoor scenes.
3) Extensive experiments on real-world and synthetic datasets show that DGFDNet achieves state-of-the-art dehazing performance with competitive computational efficiency.
II Related Work
Prior-based dehazing methods. Early approaches employ handcrafted priors to constrain the solution space, including haze-lines [20], color ellipsoidal priors [8], color-lines [21], dark channel (DCP) [5], and color attenuation [6]. Among these, DCP is the most widely used, assuming that in haze-free images, at least one color channel is nearly zero in most local regions, while haze increases intensity, enabling haze distribution estimation. However, these priors struggle in complex scenes due to discrepancies with real-world conditions.
CNNs-based dehazing methods. CNNs automatically learn features from large datasets, offering better adaptability than priors. Early models such as DehazeNet [9], AOD-Net [22], and MSCNN [23] laid the foundation for deep learning in dehazing. Later approaches, like GridDehazeNet [24] and FFA-Net [25], integrate attention mechanisms to enhance feature extraction. MSAFF-Net [26] further enhances multi-scale learning with spatial attention and feature fusion, while more recent approaches like DFR-Net [27] and DEA-Net [28] focus on fine detail recovery through task-specific modules. However, CNNs struggle with long-range dependencies, and increasing depth introduces higher computational costs.
Transformer-based dehazing methods. Due to strong global modeling capabilities, Transformers [29, 30, 31, 32, 33] are popular in image restoration, including dehazing. Dehamer [34] first integrates Transformer, refining local CNN features with global representations. DehazeFormer [14] adapts the Swin Transformer for dehazing, while STH [35] employs a multi-branch structure to process haze information at different intensities. DehazeDCT [36] incorporates deformable convolutions to handle non-uniform haze. MB-TaylorFormer V2 [37] introduces a Taylor series-inspired approximation to reduce Transformer complexity. Despite these advances, Transformer-based methods still suffer from quadratic complexity as resolution increases and struggle to restore local details.
Figure 3: Overview of DGFDNet. (a) It includes a Dehazing Main Branch and Prior Correction Guidance Branch (PCGB), both with a three-scale symmetric design and inter-stage information exchange. The core module, DGFDBlock, integrates (b) HAFM for global context modeling and (c) MGAM for local information modeling.
Frequency-domain dehazing methods. Recent studies have explored frequency-domain techniques for dehazing. Nehete et al. [16] propose a two-stage network for amplitude and phase processing. Cui et al. [38] introduce a dual-domain selection mechanism but focus mainly on suppressing low frequencies without full global spectral modulation. Their spatial attention also struggles with precise haze localization. Yu et al. [17] introduce a dual-guided framework, but their phase reconstruction may propagate errors if amplitude features are corrupted. These methods demonstrate the potential of frequency-aware dehazing, but often rely on loosely coupled designs that lack explicit degradation alignment across domains, which limits the integration of complementary features.
III Method
As shown in Fig. 3 (a), DGFDNet consists of a dehazing main branch and a prior correction guidance branch. The main branch uses a U-shaped encoder-decoder with three symmetric stages, each containing multiple DGFDBlocks. Downsampling and upsampling use standard techniques, with skip connections transferring features between the encoder and decoder. The guidance branch leverages the dark channel prior to assist HAFM in haze perception, while receiving feedback from MGAM for dynamic prior refinement. It follows the same downsampling and upsampling structure as the main branch to ensure consistency.
Given a hazy image $I\in\mathbb{R}^{3\times H\times W}$, a $3\times 3$ convolution extracts shallow features $X_{f}\in\mathbb{R}^{C\times H\times W}$, while the dark channel prior is computed as in [5]. The shallow features and the prior are fed into the main and guidance branches, respectively, which interact bidirectionally to generate restored features. A final $3\times 3$ convolution produces a residual image $I_{r}\in\mathbb{R}^{3\times H\times W}$, and the dehazed output is $I_{c}=I+I_{r}$.
III-A Haze-Aware Frequency Modulator
III-A1 Motivation
The spectral characteristics of haze degradation (Fig. 2) suggest that frequency-domain processing is more effective than spatial-domain methods for separating haze from the background. However, global modulation in the frequency domain may reduce sensitivity to subtle degradations, leading to distortion of fine details. To address this, many methods [16, 17] adopt dual-branch architectures that process spatial and frequency domains separately before merging them. However, such loosely coupled designs often limit the exchange of complementary information.
In contrast, our HAFM alternates between spatial and frequency-domain processing to maximize synergy. The spatial domain first identifies haze-affected regions, providing clear guidance for frequency-domain restoration. Specifically, the dark channel prior generates a spatial attention map that localizes degraded areas and quantifies their severity. This attention mechanism inherently enhances haze-related frequency bands, allowing frequency-domain processing to efficiently capture haze features and apply precise modulation. This explicit alignment of degradation cues enhances global context modeling while enabling fine-grained degradation perception for comprehensive dehazing.
III-A2 Pipeline
For the $j$-th DGFDBlock at the $i$-th stage, given the input feature $X_{f}^{i,j-1}$ and the dark channel-guided feature $X_{d}^{i,j-1}$, HAFM processes them as:
$\hat{X}_{f}^{i,j},\hat{X}_{d}^{i,j}=X_{f}^{i,j-1}+\mathrm{HAFM}(\mathrm{BN}(X_{f}^{i,j-1}),X_{d}^{i,j-1}),$ (1)
where $\mathrm{BN}$ represents BatchNorm, $\hat{X}_{f}^{i,j}$ is the residual output, and $\hat{X}_{d}^{i,j}$ is the intermediate dark channel-guided feature.
Spatial Modulation. As shown in Fig. 3 (b), the dark channel-guided feature $X_{d}^{i,j-1}$ is passed through two $1\times 1$ convolutions with GELU activation, producing the intermediate feature $\hat{X}_{d}^{i,j}$, which is then processed by a Sigmoid function to produce the spatial attention map $M_{sa}$:
$\hat{X}_{d}^{i,j}=\mathrm{Conv}_{1\times 1}(\mathrm{GELU}(\mathrm{Conv}_{1\times 1}(X_{d}^{i,j-1}))),$ (2)
$M_{sa}=\mathrm{Sigmoid}(\hat{X}_{d}^{i,j}).$ (3)
Meanwhile, the input feature $\mathrm{BN}(X_{f}^{i,j-1})$ undergoes a $1\times 1$ convolution to generate $X_{m}^{i,j}$ for inter-channel interaction. A $3\times 3$ depth-wise convolution is used to enhance high-frequency areas. The result is element-wise multiplied by $M_{sa}$ and added to $X_{m}^{i,j}$ to produce the spatially modulated feature:
$X_{spatial}^{i,j}=\mathrm{DConv}_{3\times 3}(X_{m}^{i,j})\odot M_{sa}+X_{m}^{i,j}.$ (4)
Figure 4: The detailed structure of PCGB. PCGB follows the three-scale design of the dehazing main branch, progressively fusing original and feedback dark channel features through SKFusion, which is shared across all DGFDBlocks at each stage.
Frequency Modulation. The spatially modulated feature $X_{spatial}^{i,j}$ is then subjected to frequency modulation. First, the feature undergoes Fast Fourier Transform (FFT, $\mathcal{F}$) to extract its real and imaginary components, $X_{\mathcal{R}}^{i,j}$ and $X_{\mathcal{I}}^{i,j}$, which are processed separately using two $1\times 1$ convolutions with GELU activation. The modulated components are subsequently converted back to the spatial domain via inverse FFT (IFFT, $\mathcal{F}^{-1}$), resulting in the frequency-modulated feature $X_{frequency}^{i,j}$:
$X_{\mathcal{R}}^{i,j},X_{\mathcal{I}}^{i,j}=\mathcal{F}(X_{spatial}^{i,j}),$ (5)
$\hat{X}_{\mathcal{R}}^{i,j},\hat{X}_{\mathcal{I}}^{i,j}=\mathrm{Conv}_{1\times 1}(\mathrm{GELU}(\mathrm{Conv}_{1\times 1}(X_{\mathcal{R}}^{i,j},\hat{X}_{\mathcal{I}}^{i,j}))),$ (6)
$X_{frequency}^{i,j}=\mathcal{F}^{-1}(\hat{X}_{\mathcal{R}}^{i,j},\hat{X}_{\mathcal{I}}^{i,j}).$ (7)
Finally, the spatially and frequency-modulated features are concatenated and fused using a $1\times 1$ convolution and GELU activation to produce the final output:
$\hat{X}_{f}^{i,j}=\mathrm{GELU}(\mathrm{Conv}_{1\times 1}(\mathrm{Cat}(X_{spatial}^{i,j},X_{frequency}^{i,j}))).$ (8)
III-B Multi-level Gating Aggregation Module
III-B1 Motivation
While HAFM effectively captures global degradation, it has limited capacity for modeling fine-detail, which is crucial for high-quality restoration. To remedy this, we propose MGAM to refine HAFM-modulated features and enhance local representation.
Recent methods [33, 18] focus on multi-scale structures to handle non-uniform haze by capturing spatial patterns at various scales. MGAM achieves this by stacking small convolutions that efficiently expand the receptive field while maintaining computational efficiency, in contrast to the large-kernel convolutions used in DSANet [39].
In addition, MGAM incorporates a hybrid gating mechanism that uses low-level features, rich in edges and textures, to regulate the flow of high-level semantic information. This hierarchical interaction enables adaptive feature fusion while preserving fine details. The sensitivity of low-level features to local haze variations further improves semantic consistency and robustness under non-uniform haze conditions.
Furthermore, the refined features from MGAM adjust the haze confidence map of the current DGFDBlock, which is then fed back to the PCGB to adaptively update the dark channel-guided features, enhancing haze localization.
III-B2 Pipeline
Given the output feature $\hat{X}_{f}^{i,j}$ from HAFM and the intermediate dark channel-guided feature $\hat{X}_{d}^{i,j}$, MGAM processes them as follows:
$X_{f}^{i,j},\tilde{X}_{d}^{i,j}=\hat{X}_{f}^{i,j}+\mathrm{MGAM}(\mathrm{BN}(\hat{X}_{f}^{i,j}),\hat{X}_{d}^{i,j}),$ (9)
where $X_{f}^{i,j}$ is the residual output of MGAM, and $\tilde{X}_{d}^{i,j}$ is the dark channel correction feature.
Multi-scale Feature Generation. As shown in Fig. 3 (c), MGAM consists of two parallel branches with gating mechanisms operating at different scales. The input feature (after BN) is passed through a $1\times 1$ convolution to double the channel count, then split into two parts: one for feature extraction ($X_{fea}^{i,j}$) and the other for gating signals ($X_{gate}^{i,j}$):
$X_{fea}^{i,j},\ X_{gate}^{i,j}=\mathrm{Split}(\mathrm{Conv}_{1\times 1}(\mathrm{BN}(\hat{X}_{f}^{i,j}))).$ (10)
Each branch applies two depth-wise convolutions for multi-scale feature extraction. The first convolution uses $k\times k$ kernels ($k\in{3,5}$), followed by a dilated convolution (dilation factor of 2) with the same kernel size to expand the receptive field, producing multi-scale features $X_{fea_{k}}^{i,j}$. Meanwhile, gating signals $X_{gate_{k}}^{i,j}$ are generated via depth-wise convolution with $k\times k$ kernels and a Sigmoid activation. The gated multi-scale features $X_{gated_{k}}^{i,j}$ are obtained by element-wise multiplication:
$X_{fea_{k}}^{i,j}=\mathrm{DDConv}_{k\times k}(\mathrm{DConv}_{k\times k}(X_{fea}^{i,j})),$ (11)
$X_{gate_{k}}^{i,j}=\mathrm{Sigmoid}(\mathrm{DConv}_{k\times k}(X_{gate}^{i,j})),$ (12)
$X_{gated_{k}}^{i,j}=X_{fea_{k}}^{i,j}\odot X_{gate_{k}}^{i,j}.$ (13)
Multi-scale Feature Fusion. The gated features are concatenated and passed through a $1\times 1$ convolution with GELU activation to produce the fused feature $X_{dual}^{i,j}$. To mitigate early-stage gating instability, a skip connection with the original input is added before a final $1\times 1$ convolution and GELU activation, yielding the multi-scale fusion feature $X_{mult}^{i,j}$:
$X_{dual}^{i,j}=\mathrm{GELU}(\mathrm{Conv}_{1\times 1}(\mathrm{Cat}(X_{gated_{3}}^{i,j},X_{gated_{5}}^{i,j}))),$ (14)
$X_{mult}^{i,j}=\mathrm{GELU}(\mathrm{Conv}_{1\times 1}(\mathrm{Cat}(\mathrm{BN}(\hat{X}_{f}^{i,j}),X_{dual}^{i,j}))).$ (15)
Feature Feedback Module. At the end of MGAM, a CAM [40] is used to generate feedback for PCGB, helping reduce redundancy in multi-scale feature extraction. The attention map $M_{ca}$ is computed from $X_{mult}^{i,j}$ and applied element-wise multiplication to both $X_{mult}^{i,j}$ and $\hat{X}_{d}^{i,j}$:
$X_{f}^{i,j}=\mathrm{Conv}_{1\times 1}(X_{mult}^{i,j}\odot M_{ca}),$ (16)
$\tilde{X}_{d}^{i,j}=\mathrm{Conv}_{1\times 1}(\hat{X}_{d}^{i,j}\odot M_{ca}).$ (17)
Here, $X_{f}^{i,j}$ is the final output, and $\tilde{X}_{d}^{i,j}$ is the feedback correction of dark channel feature. This operation allows PCGB to refine dark channel features based on channel information from $X_{mult}^{i,j}$, enhancing its alignment with the dehazing branch.
III-C Prior Correction Guidance Branch
III-C1 Motivation
The dark channel prior often fails in outdoor scenes with sky regions, where naturally low intensities lead to inaccurate transmission estimation. Moreover, its assumption of locally uniform haze does not hold in real-world conditions with complex backgrounds and varying haze densities, limiting its accuracy in haze localization. Despite these limitations, physically grounded priors offer useful inductive bias, allowing better generalization to unseen conditions. For instance, Dehamer [34] integrates the dark channel prior into positional encodings to enhance haze perception, but it treats the prior as static and lacks correction during inference, which can lead to suboptimal guidance when the prior is inaccurate.
To resolve this, we propose the Prior Correction Guidance Branch (PCGB). Instead of using the dark channel prior as a fixed input, PCGB treats it as a dynamic haze-awareness cue, refining guidance features through feedback from the dehazing branch. This preserves the inductive bias benefits while enhancing haze localization in complex, uneven conditions. Furthermore, each refinement output is fused with the initial dark channel features using SKFusion [14], combining self-correction and supervision to prevent error accumulation.
III-C2 Pipeline
As illustrated in Fig. 4, given the dark channel $X_{d}\in\mathbb{R}^{H\times W}$ of the hazy image $I$, we first encode the features using two successive convolutions:
$X_{d}^{1}=\mathrm{GELU}(\mathrm{Conv}_{1\times 1}(\mathrm{Conv}_{5\times 5}(X_{d}))),$ (18)
where $X_{d}^{1}\in\mathbb{R}^{C\times H\times W}$ is the initial dark channel feature for the first stage. Downsampling is then applied to generate the initial features for the second and third stages:
$X_{d}^{i}=f_{dw}(X_{d}^{i-1}),\quad i\in\{2,3\}.$ (19)
Dark Channel-Guided Features in the Encoder. In the first stage, the first DGFDBlock uses $X_{d}^{1}$ as the dark channel-guided feature, denoted as $X_{d}^{1,0}=X_{d}^{1}$. For subsequent stages, the first DGFDBlock receives its dark channel-guided feature by downsampling the feedback correction from the last DGFDBlock of the previous stage, and fusing it with the initial dark channel feature of the current stage via SKFusion. For the remaining DGFDBlocks in each stage, the dark channel-guided feature is obtained by fusing the feedback correction from the previous DGFDBlock with the initial dark channel feature of the current stage. The process is defined as:
$X_{d}^{i,j-1}=\begin{cases}X_{d}^{1},&i=1,j=1\\ \mathrm{SKF}_{i}(\tilde{X}_{d}^{i,j-1},X_{d}^{i}),&1<j\leq N_{i}\\ \mathrm{SKF}_{i}(f_{dw}(\tilde{X}_{d}^{i-1,N_{i-1}}),X_{d}^{i}).&i>1,j=1\end{cases}$ (20)
Here, $X_{d}^{i,j-1}$ represents the dark channel-guided feature for the $j$-th DGFDBlock at the $i$-th stage. $N_{i-1}$ is the number of DGFDBlocks in the previous stage, and $\tilde{X}_{d}^{i,j-1}$ denotes the feedback correction of dark channel from the previous DGFDBlock. $\mathrm{SKF}i$ is the SKFusion block shared across all DGFDBlocks within the $i$-th stage.
Dark Channel-Guided Features in the Decoder. The decoder follows the reverse process, propagating from Stage 3 to Stage 1. Unlike the encoder, the first DGFDBlock in each stage derives its dark channel-guided feature by upsampling the feedback correction of dark channel from the last DGFDBlock of the previous stage and fusing it with the initial dark channel feature of the current stage using SKFusion:
[…]
Here, $X_{d}^{i,j-1}$ is the dark channel-guided feature for the $j$-th DGFDBlock at the $i$-th stage in the decoder. $N_{i+1}$ denotes the number of DGFDBlocks in the next stage.
TABLE I: Quantitative comparisons with state-of-the-art dehazing methods on the synthetic and real-world datasets.
Method Venue & Year SOTS-Indoor SOTS-Outdoor Dense-Haze NH-HAZE Overhead
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM Params FLOPs
GridDehazeNet [24] ICCV 2019 32.16 0.984 30.86 0.982 13.31 0.37 13.80 0.54 0.956M 21.49G
MSBDN [41] CVPR 2020 33.67 0.985 33.48 0.982 15.37 0.49 19.23 0.71 31.35M 41.54G
FFA-Net [25] AAAI 2020 36.39 0.989 33.57 0.984 14.39 0.45 19.87 0.69 4.456M 287.8G
PMNet [42] ECCV 2022 38.41 0.990 34.74 0.985 16.79 0.51 20.42 0.73 18.90M 81.13G
MAXIM-2S [43] ICCV 2022 38.11 0.991 34.19 0.985 - - - - 14.10M 216.0G
Dehamer [34] CVPR 2022 36.63 0.988 35.18 0.986 16.62 0.56 20.66 0.68 132.50M 60.3G
Fourmer [44] ICML 2023 37.32 0.990 - - 15.95 0.49 - - 1.29M 20.6G
DehazeFormer [14] TIP 2023 40.05 0.996 34.29 0.983 - - 19.11 0.66 4.634M 48.64G
FocalNet [38] ICCV 2023 40.82 0.996 37.71 0.995 17.07 0.63 20.43 0.79 3.74M 30.63G
DEA-Net [28] TIP 2024 40.20 0.993 36.03 0.989 - - - - 3.65M 34.19G
OKNet [45] AAAI 2024 40.79 0.996 37.68 0.995 16.92 0.64 20.48 0.80 4.72M 39.71G
DCMPNet [46] CVPR 2024 42.18 0.997 36.56 0.993 - - - - 17.36M 69.23G
PGH2Net [47] AAAI 2025 41.70 0.996 37.52 0.989 17.02 0.61 - - 1.76M 16.05G
DGFDNet (Ours) - 42.18 0.997 38.51 0.995 18.34 0.67 20.49 0.81 2.08M 13.65G
III-D Loss Function
To enhance both fine-grained texture restoration and global appearance consistency, we adopt a dual-domain $L_{1}$ loss that supervises both the spatial and frequency domains. The total loss is defined as:
$\mathcal{L}=\|I_{c}-I_{g}\|_{1}+\lambda\|\mathcal{F}(I_{c})-\mathcal{F}(I_{g})\|_{1},$ (22)
where $I_{c}$ is the dehazed output of DGFDNet, $I_{g}$ is the ground-truth image, and $\lambda$ is set to $0.1$ to control the balance between the two domain losses, following the setting in [38, 45].
IV Experiments
IV-A Experimental settings
Datasets. We evaluate DGFDNet on synthetic and real-world datasets. The synthetic RESIDE dataset [48] has two training sets: ITS with 13,990 indoor pairs and OTS with 313,950 outdoor pairs. The test set, SOTS, includes 500 indoor and 500 outdoor pairs. For real-world datasets, we use Dense-Haze [49] and NH-HAZE [50], each with 55 pairs. The last five pairs from each are used for testing, the rest for training.
Comparison settings. We compare DGFDNet with ten CNN-based methods: GridDehazeNet [24], MSBDN [41], FFA-Net [25], PMNet [42], MAXIM-2S [43], DEA-Net [28], FocalNet [38], OKNet [45], PGH2Net [47], and DCMPNet [46], along with three Transformer-based models: DeHamer [34], DehazeFormer [14], and Fourmer [44]. Performance is evaluated using PSNR and SSIM.
Training details. Each stage of the model consists of {2, 2, 4, 2, 2} DGFDBlocks with a base channel size of 32. We use the Adam optimizer ($\beta_{1}=0.9,\beta_{2}=0.999$) with a dataset-dependent initial learning rate, decaying to $1\times 10^{-6}$ via cosine annealing. FLOPs are calculated for $256\times 256$ patches. All experiments run on an NVIDIA 3090 GPU.
IV-B Comparison with State-of-the-art Methods
IV-B1 Quantitative Comparisons
Table I quantitatively compares DGFDNet with SOTA methods on two synthetic and two real-world datasets. Bold and underlined values indicate the best and second-best results, respectively. DGFDNet outperforms all existing methods across all datasets.
Compared to traditional CNN-based methods like FocalNet [38] and OKNet [45], DGFDNet improves PSNR by 1.36 dB and 1.39 dB on SOTS-Indoor, and by 0.80 dB and 0.83 dB on SOTS-Outdoor, while using about half the parameters and less than half the FLOPs. Unlike FocalNet, which uses dual-domain selection with spatial attention, DGFDNet incorporates dark channel-guided attention for more accurate haze localization and a full-frequency modulation unit to better handle varying degradation levels.
Compared to the heavy DCMPNet [46], DGFDNet achieves comparable results on SOTS-Indoor with just 12.0% of the parameters and 19.7% of the FLOPs. On SOTS-Outdoor, it outperforms DCMPNet by 1.95 dB in PSNR and 0.002 in SSIM, showing better robustness in complex conditions. This improvement is driven by the PCGB module, which mitigates the limitations of the original dark channel prior. With similar resource budgets, DGFDNet consistently surpasses the lightweight PGH2Net [47], with 0.48 dB higher PSNR on SOTS-Indoor and 0.99 dB PSNR and 0.006 SSIM improvements on SOTS-Outdoor, balancing accuracy and efficiency.
On real-world datasets like Dense-Haze [49] and NH-HAZE [50], DGFDNet performs excellently. While its PSNR on NH-HAZE is slightly lower than Dehamer [34], it ranks second overall and achieves the highest SSIM on both datasets, demonstrating strong generalization and excellent global structure reconstruction in real-world hazy conditions.
IV-B2 Qualitative Comparisons
Fig. 5 shows dehazing results on synthetic images from the SOTS dataset. Our method achieves the highest PSNR across all test images. In the first row (b, c), FFA-Net [25] and MAXIM-2S [43] produce blurred edges and artifacts in complex indoor scenes. MAXIM-2S [43] and Dehamer [34] also struggle with haze removal in deeper regions (2nd row, c, d). While FocalNet [38] and OKNet [45] perform better, our method still outperforms them. As shown in Fig. 5 (f), our approach removes haze more effectively and preserves sharp edges, achieving superior visual quality. In outdoor scenes, our method continues to outperform others, achieving better haze removal and clearer reconstruction.
Fig. 6 and Fig. 7 show real-world dehazing results on the Dense-Haze [49] and NH-HAZE [50] datasets. In Fig. 6, two dense haze scenes are compared. In the first, our method achieves slightly lower PSNR than Dehamer [34], but surpasses it by 0.183 in SSIM, offering clearer textures and reduced color distortion. In the second scene, with severe haze and significant detail loss, our method restores the overall structure and preserves more details than Dehamer [34].
Fig. 7 compares two non-homogeneous haze scenes. Apart from a slightly lower PSNR in the second scene compared to FocalNet [38], our method outperforms others in all metrics. In the first scene, it minimizes color distortion and avoids heavy artifacts from other methods. In the second scene, it removes haze most effectively, revealing clearer details, though it inadvertently eliminates some snow traces.
Overall, our method achieves the highest SSIM across all real-world cases, demonstrating strong global structure recovery even under severe haze-induced detail degradation.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
9.79 dB 32.69 dB 34.27 dB 32.18 dB 33.67 dB 32.58 dB 35.20 dB PSNR
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
11.54 dB 34.49 dB 30.01 dB 34.38 dB 40.42 dB 40.29 dB 41.06 dB PSNR
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
22.04 dB 31.72 dB 29.04 dB 28.88 dB 35.11 dB 35.82 dB 39.72 dB PSNR
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
12.35 dB 26.62 dB 28.41 dB 26.86 dB 30.09 dB 31.83 dB 33.39 dB PSNR
(a) Hazy image (b) FFA-Net [25] (c) MAXIM-2S [43] (d) Dehamer [34] (e) FocalNet [38] (f) OKNet [45] (g) Ours (h) GT
Figure 5: Visual comparisons on synthetic hazy images from the SOTS dataset. Key regions highlighted by red boxes are enlarged in the lower-left corner for clearer comparison.
TABLE II: Ablation study for three core modules.
Variant SOTS-Indoor Overhead
PSNR SSIM Params FLOPs
MGAM 40.10 0.994 0.88M 8.39G
HAFM 39.35 0.994 1.12M 4.69G
HAFM+PCGB 40.05 0.995 1.48M 7.26G
MGAM+HAFM 41.90 0.996 2.03M 11.94G
Full Network 42.18 0.997 2.08M 13.65G
IV-C Ablation Studies
We first evaluate the individual and combined contributions of HAFM, MGAM, and PCGB to verify their effectiveness and complementarity. Ablation studies are then conducted by modifying each module independently while keeping the rest of the architecture fixed. All models are trained on the ITS dataset and evaluated on SOTS-Indoor under the same conditions as the final model. Since PCGB targets the limitations of dark channel priors in outdoor scenes, we also validate its performance on SOTS-Outdoor using a model trained on the OTS dataset. Finally, attention map visualizations are provided to demonstrate the corrective effect of PCGB.
TABLE III: Ablation study for key components in HAFM.
Variant SOTS-Indoor Overhead
PSNR SSIM Params FLOPs
HAFM-S 41.12 0.996 1.31M 12.78G
HAFM-F 41.18 0.996 1.83M 10.95G
HAFM-SSA 41.78 0.996 1.84M 11.13G
HAFM-SMLP 41.91 0.997 2.09M 20.22G
Full HAFM 42.18 0.997 2.08M 13.65G
TABLE IV: Ablation study for key components in MGAM.
Variant SOTS-Indoor Overhead
PSNR SSIM Params FLOPs
MGAM-$3\times 3$ 41.79 0.995 1.93M 11.74G
MGAM-$5\times 5$ 41.58 0.996 1.97M 12.45G
MGAM-Nogate 41.15 0.996 1.97M 12.34G
MGAM-Noskip 41.76 0.996 1.91M 12.01G
Full MGAM 42.18 0.997 2.08M 13.65G
IV-C1 Effectiveness of Core Modules
To thoroughly evaluate the effectiveness and complementarity of the three core modules, we design four variants: MGAM, HAFM (without feedback correction), HAFM+PCGB (with full feedback-guided correction), and MGAM+HAFM. Table II summarizes the performance of each variant.
As shown in Table II, MGAM improves PSNR by 0.75 dB over HAFM while maintaining the same SSIM of 0.994, highlighting its advantage in enhancing local details. Introducing PCGB into HAFM yields an additional 0.70 dB gain, emphasizing its role in enhancing global degradation modeling via adaptive prior correction. Combining MGAM and HAFM results in a 1.80 dB PSNR improvement over MGAM alone and 2.55 dB over HAFM, benefiting from the synergy between multi-scale spatial feature extraction and dual-domain modulation. Further adding PCGB achieves the best overall performance, confirming that the three modules complement each other and jointly contribute to both local and global restoration with minimal additional computational cost.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Hazy image (b) Dehamer [34] (c) Ours (d) GT
Figure 6: Visual comparisons on real-world hazy images from the Dense-Haze dataset. Key regions marked with red boxes are enlarged in left-to-right order and arranged vertically on the right.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Hazy image (b) Dehamer [34] (c) FocalNet [38] (d) Ours (e) GT
Figure 7: Visual comparisons on real-world hazy images from the NH-Haze dataset. Key regions marked with red boxes are enlarged in left-to-right order and arranged horizontally at the bottom.
IV-C2 Effectiveness of HAFM
To assess the role of HAFM, we design four variants: HAFM-S (spatial modulation only), HAFM-F (frequency modulation only), HAFM-SSA (standard spatial attention replacing dark channel-guided attention), and HAFM-SMLP (spatial MLP replacing frequency modulation). The results are presented in Table III.
According to Table III, using only spatial (HAFM-S) or only frequency modulation (HAFM-F) results in PSNR drops of 1.06 dB and 1.00 dB, respectively, compared to the full HAFM. This underscores the importance of combining both spatial and frequency modulation for optimal performance. Replacing dark channel-guided attention with standard spatial attention (HAFM-SSA) leads to a 0.40 dB drop in PSNR and a 0.001 decrease in SSIM, highlighting that the haze-specific spatial modulation offers a stronger prior to frequency modulation. Substituting the frequency-domain MLP with a spatial MLP (HAFM-SMLP) increases FLOPs by 48.1%, yet PSNR remains 0.27 dB lower than the full HAFM, showing that frequency-domain processing not only boosts computational efficiency but also enhances global feature modeling.
IV-C3 Effectiveness of MGAM
To assess the contribution of each MGAM component, we conduct an ablation study with four variants: MGAM-$3\times 3$ (only the $3\times 3$ gated branch), MGAM-$5\times 5$ (only the $5\times 5$ gated branch), MGAM-Nogate (both branches without gating), and MGAM-Noskip (removing the skip connection). The results are reported in Table IV.
TABLE V: Ablation study for key components in PCGB on SOTS-Indoor.
Variant SOTS-Indoor Overhead
PSNR SSIM Params FLOPs
PCGB-SSA 41.78 0.996 1.84M 11.13G
PCGB-NoFR 41.90 0.996 2.03M 11.94G
PCGB-DAF 42.02 0.996 2.07M 13.62G
PCGB-PFF 41.47 0.996 2.08M 13.05G
Full PCGB 42.18 0.997 2.08M 13.65G
TABLE VI: Ablation study for key components in PCGB on SOTS-Outdoor.
Variant SOTS-Outdoor Overhead
PSNR SSIM Params FLOPs
PCGB-SSA 37.81 0.995 1.84M 11.13G
PCGB-NoFR 37.53 0.994 2.03M 11.94G
PCGB-DAF 37.95 0.995 2.07M 13.62G
PCGB-PFF 37.82 0.995 2.08M 13.05G
Full PCGB 38.51 0.995 2.08M 13.65G
MGAM-$3\times 3$ achieves 0.21 dB higher PSNR but 0.001 lower SSIM compared to MGAM-$5\times 5$, indicating that small-kernel convolutions are better at capturing local textures, while large-kernel convolutions capture global structures. The full MGAM, combining both, balances these strengths, improving both PSNR and SSIM. Removing the gating mechanism (MGAM-Nogate) results in a 1.03 dB drop in PSNR and 0.001 in SSIM, emphasizing the importance of dynamic gating for adaptive feature selection and mitigating detail loss. Finally, removing the skip connection (MGAM-Noskip) reduces PSNR by 0.42 dB and SSIM by 0.001, underscoring its role in correcting selection bias from the gating mechanism.
IV-C4 Effectiveness of PCGB
To validate the rationale behind PCGB, we test four variants: (1) PCGB-SSA, which employs standard spatial attention, identical to HAFM-SSA; (2) PCGB-NoFR, where the spatial attention map is generated from the initial dark channel features without feedback refinement (3) PCGB-DAF, which directly adds feedback correction to the initial dark channel features, bypassing SKFusion [14]; (4) PCGB-PFF, which uses the initial prior only in the first DGFDBlock, while subsequent blocks fuse the guidance and feedback from the previous DGFDBlock via SKFusion. Results on the SOTS-Indoor and SOTS-Outdoor datasets are exhibited in Table V and Table VI, respectively.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Hazy image (b) GT (c) Dark channel (d) First HAFM Attnmap (e) Last HAFM Attnmap
Figure 8: Visual study of PCGB for dehazing. From left to right: hazy images, ground truth images, dark channels of the hazy images, and dark channel-guided spatial attention maps generated by the first and last HAFM.
Indoor results. Table V shows that the dark channel prior remains a reliable guidance cue in indoor scenes. Even without feedback refinement, PCGB-NoFR outperforms PCGB-SSA by 0.12 dB in PSNR, demonstrating the prior’s effectiveness in haze localization and degradation estimation. Adding feedback correction (PCGB-DAF) improves PSNR by 0.12 dB, and using SKFusion in the full PCGB increases both PSNR by 0.16 dB and SSIM by 0.001, demonstrating better adaptability to complex structures. However, PCGB-PFF, which discards the prior after the first block, results in a 0.71 dB PSNR drop, highlighting the importance of consistently using the initial prior for stable and accurate guidance.
Outdoor results. In outdoor environments, the dark channel prior becomes less reliable due to factors like sky and bright regions. As shown in Table VI, PCGB-NoFR performs 0.28 dB worse than PCGB-SSA, highlighting the limitations of the prior alone. Introducing feedback correction (PCGB-DAF) improves performance, yielding gains of 0.42 dB over PCGB-NoFR and 0.14 dB over PCGB-SSA. PCGB-PFF, which progressively fuses the prior with feedback, improves over NoFR by 0.29 dB and matches PCGB-SSA, proving the benefit of self-correction on noisy priors in outdoor scenes. However, PCGB-DAF still outperforms PCGB-PFF by 0.13 dB, emphasizing the importance of retaining the initial prior to avoid accumulating errors. The full PCGB achieves the best, surpassing PCGB-DAF by 0.56 dB, validating SKFusion’s ability to handle complex outdoor haze patterns.
IV-C5 Visual Analysis of PCGB
To further evaluate the corrective effect of PCGB, we visualize the dark channel-guided spatial attention maps from the first and last HAFM modules of DGFDNet. The visualization covers three representative scenes from the SOTS-Indoor, SOTS-Outdoor, and NH-HAZE datasets, as shown in Fig. 8.
In the indoor scene, the initial attention map follows the raw dark channel distribution, causing widespread activation. After feedback refinement, attention is focused on heavily degraded areas, improving haze localization and spatial modulation. In the outdoor scene, the limitations of dark channel priors are clear. In bright regions like the sky, the initial map overemphasizes these areas, neglecting distant buildings and roads. Feedback refinement shifts attention to the other haze regions, reducing over-processing of the sky.
In the NH-HAZE example, the initial attention map incorrectly activates on irrelevant structures, like the swing’s mesh, causing false positives. It also focuses on dense haze areas while missing subtly degraded regions. After refinement, the final map better aligns with the haze distribution, suppressing false activations and improving sensitivity to neglected regions. These results demonstrate the robustness of PCGB in handling complex, spatially diverse degradation patterns.
V Conclusion
In this paper, we propose DGFDNet, a dual-domain dehazing framework that explicitly aligns spatial and frequency-domain degradation cues under dark channel guidance. To address inefficiencies and weak coupling in existing spatial-frequency models, DGFDNet introduces three key modules: the Haze-Aware Frequency Modulator (HAFM) for degradation-aware spectral filtering, the Multi-level Gating Aggregation Module (MGAM) for adaptive multi-scale spatial fusion, and the Prior Correction Guidance Branch (PCGB) for iterative dark channel prior refinement. These modules form a compact, efficient framework that captures long-range dependencies and preserves fine details, handling both homogeneous and non-homogeneous haze. Extensive experiments on synthetic and real-world benchmarks show DGFDNet’s superior performance, strong generalization, and computational efficiency, making it well-suited for practical deployment.
Acknowledgments
This work was partially supported by National Natural Science Foundation of China (Nos.62471317), Natural Science Foundation of Shenzhen (No. JCYJ20240813141331042), and Guangdong Provincial Key Laboratory (Grant 2023B1212060076).